OpenStudy (anonymous):

-56y^7 divided by -7y^5= 8y^2, right?

OpenStudy (anonymous):

yea u r right

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

since : \[-56y^7/-7y^5 = 8y^7/y^5\] as we know that \[a^b/a^c=a^{(b-c)}\] now \[8y^7/y^5=8y^{(6-5)}\] that is \[8y^2\]

OpenStudy (anonymous):

got it